If u=3t2i+et j−t k and v=sinti−costj+e2tk then, at t=0, ddt(u×v)=
1
Expert's answer
2020-09-28T19:56:31-0400

"\mathbf{We \;have-\;\;u=3t^2\;i+e^t\;j-t\;k\;\;and\;v=sin(t)\;i-cos(t)\;j+e^{2t}\;k}"


"\mathbf{To\;find- \;at\;t=0\;,\;\dfrac{d}{dt}(u \times v)= \;?}"


"\mathbf{So,\;u \times v= \begin{vmatrix}
 i & j & k\\
 3t^2 & e^t &-t\\
\sin(t) & \cos(t) & e^{2t}
\end{vmatrix}=i\;\begin{vmatrix}
 e^t & -t \\
 \cos(t) & e^{2t}
\end{vmatrix}-j\;\begin{vmatrix}
 3t^2 & -t \\
 \sin(t) & e^{2t}
\end{vmatrix}}"


"\mathbf{+k\;\begin{vmatrix}
 3t^2 & e^t \\
 \sin(t) & \cos(t)
\end{vmatrix}}"


"\mathbf{\implies u \times v=i\;(e^t.e^{2t}-(-t).\cos(t))-j(3t^2.e^{2t}-(-t).\sin(t))}"


"\mathbf{+k\;(3t^2.\cos(t)-e^t.\sin(t))}"


"\mathbf{=i\;(e^{3t}+t\cos(t))-j(3t^2e^{2t}+t\sin(t))+k(3t^2\cos(t)-e^t\sin(t))}"


"\mathbf{Now,\;\dfrac{d}{dt}(u \times v)=\dfrac{d}{dt}\bigg[i\;(e^{3t}+t\cos(t))-j(3t^2e^{2t}+t\sin(t))}"


"\mathbf{+k(3t^2\cos(t)-e^t\sin(t))\bigg]}"


"\mathbf{=i\;\bigg(\dfrac{d}{dt}(e^{3t}+t\cos(t))\bigg)-j\;\bigg(\dfrac{d}{dt}(3t^2e^{2t}+t\sin(t))\bigg)}"


"\mathbf{+k\;\bigg(\dfrac{d}{dt}(3t^2\cos(t)-e^t\sin(t))\bigg)}"


"=" "\mathbf{i\;(3e^{3t}+\cos(t)-t\sin(t))-j\;(6t^2e^{2t}+6te^{2t}+\sin(t)+t\cos(t))}"


"\mathbf{+k\;(6t\cos(t)-3t^2\sin(t)-e^t\sin(t)-e^t\cos(t))}"


"\mathbf{At\;t=0,\; \dfrac{d}{dt}(u \times v)=i\;(3+1-0)-j\;(0+0+0+0)+k\;(0-0-0-1)}"


"\mathbf{=4\;i-0\;j+(-1)\;k=4\;i-k}" ......................Ans.